# WixUp: A General Data Augmentation Framework for Wireless Perception in Tracking of Humans

### v1

@@ Abstract
Recent advancements in wireless perception technologies, including mmWave, WiFi, and acoustics, have expanded their application in human motion tracking and health monitoring. They are promising alternatives to traditional camera-based perception systems, thanks to their efficacy under diverse conditions or occlusions, and enhanced privacy. However, the integration of deep learning within this field introduces new challenges such as the need for extensive training data and poor model generalization, especially with sparse and noisy wireless point clouds. As a remedy, data augmentation is one solution well-explored in other deep learning fields, but they are not directly applicable to the unique characteristics of wireless signals. This motivates us to propose a custom data augmentation framework, WixUp, tailored for wireless perception. Moreover, we aim to make it a general framework supporting various datasets, model architectures, sensing modalities, and tasks; while previous wireless data augmentation or generative simulations do not exhibit this generalizability, only limited to certain use cases. More specifically, WixUp can reverse-transform lossy coordinates into dense range profiles using Gaussian mixture and probability tricks, making it capable of in-depth data diversity enhancement; and its mixing-based method enables unsupervised domain adaptation via self-training, allowing training of the model with no labels from new users or environments in practice. In summary, our extensive evaluation experiments show that WixUp provides consistent performance improvement across various scenarios and outperforms the baselines.
1 Introduction
@@
The advancements in wireless perception have enabled a broad spectrum of applications, ranging from human tracking and health monitoring to radar systems for self-driving and robotics. It relies on various wireless sensing modalities including mmWave, acoustics, WiFi, etc. These technologies offer several advantages over traditional camera-based perception systems. Wireless signals are not restricted by light conditions or line-of-sight occlusions, and they provide better privacy protection compared to cameras. Typically, a wireless perception system emits custom signals such as Frequency-modulated continuous wave (FMCW) from the transmitter [32]; then by analyzing the reflections recorded by the receiver, it can extract information for localization or motion tracking.
@@
The recent emergence of utilizing deep learning for wireless perception systems highly improves their capability of tracking [62, 61], but also brings forth a set of new challenges. Firstly, there is a significant demand for collecting extensive training data that requires considerable resources and effort. Secondly, the trained model often exhibits poor generalization when deployed in new scenarios, such as a new user or environment. Moreover, when the raw FMCW signal is processed into point clouds, it exhibits severe sparsity issues that affect the training data quality. Consequently, downstream tasks encounter limitations in accuracy and struggle to address complex tasks such as multi-person tracking or flow estimation. For instance, Fig. 2 demonstrates two samples of the processed mmWave as point clouds, which appear sparse and barely resemble a coherent human skeletal structure; while the depth camera provides dense point clouds. The disparity is owing to the inherent limitations of wireless technology, where its low granularity is insufficient for multi-part tracking.
@@
To address these challenges, data augmentation is an effective solution that has been proven in many fields of machine learning [34, 57]. There are well-explored methods for augmenting 2D or 3D data like images [57] and Lidar point clouds [51], broadly categorized into global augmentation and local augmentation [51]. 1) Global augmentation performs random scaling, flipping, and rotation on 2D images or 3D Lidar point clouds to transform data. However, they fail to augment local structures and ignore the relevance between elements within one data sample. 2) Instead, local augmentation techniques tend to mix data to generate new samples. The notable mixup [57] paper uses convex combination to mix data, and  [56, 51] crop&paste local structure from one data to another, at a variety of granularity such as a patch or an object. Additional benefits of mixing-based data augmentation are that it can not only improve model performance in supervised learning but also perform unsupervised domain adaptation via self-training[51], which helps with reducing labeling efforts. Nevertheless, mixing-based augmentation requires designing a domain-specific patch/object selection algorithm.
@@
In current wireless research, some work has adopted straightforward global data augmentation methods such as rotation, Gaussian noise to classify RF signal types [23, 11, 16, 46] or shifting the range or initial phase for perception [1]. However, mixing-based local data augmentation remains unexplored because of its complexity in incorporating domain-specific representations. Also, these previous methods are confined to specific tasks or models and often lack a comprehensive evaluation.
Besides, a recent trend involves leveraging generative models to simulate and synthesize wireless data [64, 14], sharing similar motivation with data augmentation; but they are also limited to certain tasks or data. For instance, [5, 59] uses video data of the human body to simulate RF signals, but it can only generate scenes from existing video data. [14] overcomes this issue by using text prompts to synthesize the 3D mesh of human actions, and then simulate mmWave data from the 3D visuals using ray tracing and Diffussion models. Nevertheless, this only supports action-related tasks, because the text prompts can only describe body actions or hand gestures. For example, it can only synthesize data for pose estimation or action recognition tasks, but not for user identification, since it is intractable to accurately depict behavior traits solely by text prompts. Besides, [64] uses NeRF to simulate RF signals to aid simple tasks of localization or 5G channel estimation, i.e. static scene synthesis other than motions; and it needs labeled data from the environment to train the NeRF first. Overall, generative methods demand non-trivial effort to train an extra model and do not generalize to out-of-distribution data.
@@
In contrast, in this work, we aim to propose a mixing-based data augmentation framework, that is training-free and particularly good at increasing data diversity or even closing domain gaps [51]. It supports multiple human tracking tasks, model architectures, and sensing modalities. Moreover, beyond improving model performance, it could also reduce labeling efforts by unsupervised domain adaptation via self-training. In practice, this enables training the model with new user/environment data without any labels. These benefits make our framework much more practical for wireless sensing than either conventional global augmentation or generative simulation.
@@
However, this goal poses two primary challenges: 1) We need a mixing algorithm tailored to the features of wireless signals. Unlike 2D images or Lidar point clouds, wireless raw data is less interpretable, rendering many conventional augmentation strategies not applicable, like flipping. Even after processing the raw data into point clouds, the noise and the sparsity issue we discussed in Fig. 2 make general 3D data augmentation insufficient. 2) To perform a comprehensive evaluation of our method on diverse real-world data and tasks, we need access to multiple high-quality datasets. Excitingly, in the realm of mmWave sensing, we have started seeing trends of open-sourcing data recently [17, 53, 7], but they do not provide the unprocessed raw data. The processed point clouds are lossy information, because of the constant false alarm rate (CFAR) filtering as part of the standard FMCW processing pipeline.
@@
Thereby, we propose WixUp, a range-profile level data augmentation tailored for the unique characteristics in wireless perception, incorporating the Gaussian mixture to solve the lossy point cloud issue. To elaborate, we employ a custom data processing pipeline to transform any format of raw data into the lossless space of range profile, by simulating irreversible transformation of CFAR using Gaussian mixture. Then the mixing is equivalent to the intersection of two Gaussian mixtures, along with probability-based methods to bootstrap the results and embody spherical angles. Finally, by framing a unified code base, we meticulously investigate the generalizability of our data augmentation framework across three datasets, two model architectures, three tasks, and two sensing modalities of mmWave and acoustics, with a focus on human perception because this topic has most open research resource available to make this study ready. Furthermore, we demonstrate its efficacy in unsupervised domain adaptation via self-training, across unseen environments and users. Experiment results show that our data augmentation framework exhibits a consistent improvement across different evaluation experiments over baselines.
In summary, our paper has the following main contributions:
@@
  • We propose a custom mixing-based data augmentation in wireless perception, which preserves characteristic representations for in-depth augmentation and handles imperfect data types.
  • Our pipeline is a general data augmentation framework across datasets, tasks, model architectures, and sensing modalities. And we conduct comprehensive experiments in a unified setup to validate its generalizability.
@@
2 Related work
@@
Data augmentation techniques (DA) is well-explored for 2D images [22, 45, 39] and 3D Lidar point clouds [51], since collecting and annotating training data is labor-intensive and time-consuming. DA of 2D images typically refers to global augmentation techniques such as random cropping [26, 44], scaling [42, 44], erasing [65], and color jittering [44], which aim to learn transformation invariance for image recognition tasks. In contrast, local augmentation techniques, such as mixup [57] and CutMix [56], generate new training data through various mixing operations. 3D DA has garnered attention recently, which also adopts similar methods such as scaling, rotation, and translation; Recent studies attempt to augment local structures of point clouds [30, 40, 25] and introduce concepts for object-level augmentation [15, 27] and scene-level [51] in self-driving. However, conventional augmentation techniques are not suitable for addressing the unique characteristics of wireless sensing data, as we discussed above, necessitating the development of customized approaches.
@@
Data augmentation for labeling efficiency. The mixing-based local DA can not only improve model performance in standard supervised learning, but also reduce labeling efforts via self-training [51]. It enables unsupervised domain adaptation without explicit supervision. Its effectiveness has been studied in text analysis, computer vision [55, 18, 51], and speech processing [41], improving the generalizability of models. It synthesizes silver standard labels generated from input data to facilitate learning. Similarly, in self-supervised learning, DA [9, 10] even brings more performance improvements than algorithmic techniques like revising model architectures or algorithmic additions [12, 13, 19, 8]. This aspect is crucial for wireless perception because, unlike crawling images and text from the web, collecting sensing data with labels from the real physical world is especially expensive.
@@
Wireless perception is a promising alternative to cameras for tracking humans, relying on the modalities of radio frequency (RF) signals, such as mmWave and WiFi, or acoustics. The applications range from gesture classification [21, 48], localization [60, 38, 37], motion detection and pose estimation [4, 3, 2, 62, 61, 24], and fine-grained face reconstruction [52] using RF. In addition, acoustics detection leverages ubiquitous speakers and microphones with little bandwidth for hand gesture recognition [47, 50, 20, 54] and hand tracking [36, 49, 33, 31, 28].
@@
Wireless perception datasets have become more accessible in recent research. Some focus primarily on a single task, with a majority being devoted to keypoint estimation or action recognition using mmWave [43, 63]. MARS [7] is one of the pioneers, providing data for rehabilitation using mmWave. mRI [6] and MM-Fi [53] are large-scale structures that contain around 160K frames. MiliPoint [17] first includes all three main tasks in human tracking: user identification, action classification, and keypoint pose estimation, with a total of 49 different actions across 545k frames. These datasets facilitate research on wireless perception and make DA research possible in this domain by being able to benchmark comprehensively.
@@
Data augmentation in wireless perception was previously mentioned in some work, typically tailored to their specific system, while there is a lack of a general data augmentation framework with comprehensive evaluations. For examples of global augmentation, MiliPoint [17] employs a DA, named stack, which involves zero-padding and random resampling. Consequently, the augmented points essentially replicate the original ones with some subset and duplication. We use this as a baseline in our experiments. [23, 11, 16, 46] use global DA like flipping, rotation, Gaussian noise, or wavelet but only to classify radio signal types, not for perception.  [1] augments data by shifting the range profile along the range axis by a small scale. However, beyond a certain augmentation scale, the shift makes data distorted, leading to a decrease in model performance. Besides, [58] uses transfer learning to improve data efficiency across environments but it requires labeled data from the target domain.
@@
Recently, generative models have been investigated to synthesize wireless data [64, 14], but they are limited to certain tasks or data. [5, 59] simulates RF signals from videos of human actions, but the generation is bound to existing video data. [14] solves this issue by using text prompts to synthesize the 3D mesh of humans first, and then simulate mmWave data from the 3D visuals using ray tracing and Diffussion models. However, the text prompts are limited to the description of actions; it does not support non-action tasks like user identification, where describing behavior traits with text is intractable. NeRF is also applied to simulate RF signals[64], but it only demonstrates simple static tasks of localization or 5G channel estimation rather than motion tracking, since the vanilla NeRF is for static synthesis; and it requires labeled data from the target scene to train the NeRF model. In general, generative methods require significant effort to train an additional model and struggle to generalize to out-of-distribution data, while mixing-based DA is training-free and excels at enhancing data diversity.
3 Background: FMCW raw data processing and formats
Figure 1: The common data processing pipeline for wireless perception.
Figure 2: The sparsity issue in wireless.
Before delving into WixUp’s mixing pipeline, we first extend the background, to reveal what is challenging and motivate our custom designs for WixUp.

### v2

# WixUp: A Generic Data Augmentation Framework for Wireless Human Tracking

@@ Abstract.
Wireless sensing technologies, leveraging ubiquitous sensors such as acoustics or mmWave, can enable various applications such as human motion and health tracking. However, the recent trend of incorporating deep learning into wireless sensing introduces new challenges, such as the need for extensive training data and poor model generalization. As a remedy, data augmentation is one solution well-explored in other fields such as computer vision; yet they are not directly applicable due to the unique characteristics of wireless signals. Hence, we propose a custom data augmentation framework, WixUp, tailored for wireless human sensing. Our goal is to build a generic data augmentation framework applicable to various tasks, models, data formats, or wireless modalities. Specifically, WixUp achieves this by a custom Gaussian mixture and probability-based transformation, making any data formats capable of an in-depth augmentation at the dense range profile level. Additionally, our mixing-based augmentation enables unsupervised domain adaptation via self-training, allowing model training with no ground truth labels from new users or environments in practice. We extensively evaluated WixUp across four datasets of two sensing modalities (mmWave, acoustics), two model architectures, and three tasks (pose estimation, identification, action recognition). WixUp provides consistent performance improvement (2.79%-84.25%) across these various scenarios and outperforms other data augmentation baselines.
Wireless sensing, Wireless perception, Data augmentation, Self-training
journalyear: 2025copyright: ccconference: The 23rd ACM Conference on Embedded Networked Sensor Systems; May 6–9, 2025; Irvine, CA, USAbooktitle: The 23rd ACM Conference on Embedded Networked Sensor Systems (SenSys ’25), May 6–9, 2025, Irvine, CA, USAdoi: 10.1145/3715014.3722084isbn: 979-8-4007-1479-5/2025/05submissionid: 704
1. Introduction
@@
Recent advancements in wireless sensing have enabled a broad spectrum of applications, ranging from human tracking and health monitoring to large radar systems for self-driving and robotics. It includes various wireless sensing modalities using ubiquitous sensors, such as mmWave, acoustics, WiFi, etc. These technologies offer several advantages over traditional camera-based perception systems. Wireless signals are not restricted by light conditions or line-of-sight occlusions, and they provide better privacy protection compared to cameras. Typically, a wireless sensing system emits custom signals, such as frequency-modulated continuous wave (FMCW) from the transmitter (lovescu2020fundamentals, ); these signals are reflected by various objects and people in the environment. Then, by analyzing the reflections recorded by the receiver, it can extract information for localization or motion tracking.
@@
The recent emergence of utilizing deep learning in wireless sensing greatly improves their tracking capability (zhao2018rf, ; zhao2018through, ), but also brings forth new challenges. Firstly, there is a significant demand for collecting extensive training data that requires considerable resources and effort. Secondly, the trained model often exhibits poor generalization when deployed in new scenarios, such as new users or environments. Moreover, when the raw wireless signal is processed into point clouds, it exhibits severe sparsity issues that affect the training data quality. Consequently, downstream tasks face limitations in accuracy and struggle to address complex tasks, such as flow estimation. For instance, Fig. 2 demonstrates two samples of the processed mmWave as point clouds, which appear sparse and barely resemble a coherent human skeletal structure; while the depth camera provides dense point clouds. The disparity is owing to the inherent limitations of wireless technology, where its low granularity is insufficient for multi-part tracking.
@@
To address these challenges, data augmentation is an effective solution, proven in many fields of machine learning (morningstar2024augmentations, ; zhang2017mixup, ). There are well-explored methods for augmenting 2D or 3D data like images (zhang2017mixup, ) and Lidar point clouds (xiao2022polarmix, ), broadly categorized into global augmentation and local augmentation (xiao2022polarmix, ). 1) Global augmentation performs random scaling, flipping, and rotation on 2D images or 3D Lidar point clouds to transform individual data. However, they fail to augment local structures and ignore the relevance between elements within individuals. 2) Instead, local augmentation techniques tend to mix two data to generate new samples. Mixup (zhang2017mixup, ) uses convex combination to mix data, theoretically proved that mixing-based data augmentation is effective for generalization;  (yun2019cutmix, ; xiao2022polarmix, ) crop&paste local structure of patch or an object from one data to another. Additional benefits of mixing-based data augmentation are that, beyond improving model performance in supervised learning, it helps perform unsupervised domain adaptation via self-training(xiao2022polarmix, ), which reduces labeling efforts. Nevertheless, mixing-based augmentation requires designing a domain-specific patch/object selection algorithm.
@@
Some current wireless research has adopted straightforward global augmentation methods, such as rotation or Gaussian noise to classify RF signal types (huang2019radiomodule, ; chen2023radiowavelet, ; clark2021training, ; blindsignal2019, ), and shifting the range or initial phase for perception (acousticdataset2023, ). However, mixing-based local augmentation remains unexplored, due to its complexity in incorporating domain-specific representations. Also, the above global methods are confined to specific tasks or models and often lack a comprehensive evaluation.
Besides, a recent trend involves leveraging generative models to simulate and synthesize wireless data, sharing similar motivation with data augmentation; but they are also limited to certain tasks or data. In brief, some cross-modality synthesis work use videos(ahuja2021vid2doppler, ; 2023synthesizemili, ) or text prompts(chen2023rfgenesis, ) to generate RF signals; but they are constrained by limited source video or inadequate details conveyed by text description. Besides, (zhao2023nerf2, ; chi2024rfdiff, ) use NeRF or Diffusion to simulate RF signals, only aiding static tasks like localization and channel estimation or simple gesture classification; intensive labeled data are also required to train the model first. Overall, generative methods demand non-trivial effort to train an extra model and can hardly generate out-of-distribution data.
@@
In this work, we aim to propose a mixing-based data augmentation framework, that is training-free and particularly good at increasing data diversity or even closing domain gaps via self-training (xiao2022polarmix, ). It could support multiple human tracking tasks, model architectures, and wireless sensing modalities (like RF or acoustics). Furthermore, beyond improving model performance, it could also reduce labeling efforts by unsupervised domain adaptation via self-training. In practice, this enables training the model with new user/environment data without any labels. These benefits make our framework significantly more practical for wireless sensing than either conventional global augmentation or generative simulation. Ablation studies also justify the effectiveness and rationality of the approach.
@@
However, this goal poses two primary challenges: 1) We need a mixing algorithm tailored to the features of wireless signals. Unlike 2D images or Lidar point clouds, wireless raw data is less interpretable, rendering many conventional augmentation strategies not applicable, like flipping. Even after processing the raw data into point clouds, the noise and the sparsity issue we discussed in Fig. 2 make existing 3D data augmentations insufficient. 2) To perform a comprehensive evaluation of our method on diverse real-world data and tasks, we need access to multiple high-quality datasets. Excitingly, in the realm of mmWave sensing, we have started seeing trends of open-sourcing data recently (cui2024milipoint, ; yang2024mm, ; marsdataset2021, ), but they do not provide the unprocessed raw data. The processed point clouds are lossy information, because of the constant false alarm rate (CFAR) filtering as part of the standard FMCW processing pipeline.
@@
Thereby, we propose WixUp, a range-profile-level data augmentation tailored for the unique characteristics in wireless perception, incorporating the Gaussian mixture to solve the lossy point cloud issue. 1) To elaborate, we employ a custom data processing pipeline to transform any format of raw data into the lossless space of range profile, involving simulating irreversible transformation of CFAR. This is based on representing range profile as a Gaussian mixture of various ranges. Therefore, mixing two data points for data augmentation is equivalent to the intersection of two Gaussian mixtures. Additionally, probability-based methods can further bootstrap the augmented data; embodying spherical angles over ranges can yield 3D outputs. 2) Finally, by framing a unified code base, we meticulously investigate the generalizability of our data augmentation framework across four datasets involving two sensing modalities (mmWave and acoustics), two model architectures, and three tasks, with a focus on human perception because this topic has most open research resource available to make this study ready. Furthermore, we demonstrate its efficacy in unsupervised domain adaptation via self-training, across unseen environments and users. Experiment results show that our data augmentation framework exhibits a consistent improvement across different evaluation experiments over baselines.
In summary, our main contributions are:
@@
  • We propose a custom mixing-based (local) data augmentation technique for wireless sensing that preserves characteristic representations for in-depth augmentation and handles imperfect data types.
@@
  • We conduct comprehensive experiments in a unified setup to demonstrate WixUp’s generalizability across tasks, model architectures, datasets, and sensing modalities.
2. Related work
@@
Data augmentation techniques (DA) is well-explored for 2D images (he2016deep, ; tan2019efficientnet, ; ren2015faster, ) and 3D Lidar point clouds (xiao2022polarmix, ), since collecting and annotating training data is labor-intensive and time-consuming. DA of 2D images typically refers to global augmentation techniques such as random cropping (krizhevsky2012imagenet, ; szegedy2015going, ), scaling (simonyan2014very, ; szegedy2015going, ), erasing (zhong2020random, ), and color jittering (szegedy2015going, ), which aim to learn transformation invariance for image recognition tasks. In contrast, local augmentation techniques, such as mixup (zhang2017mixup, ) and CutMix (yun2019cutmix, ), generate new training data through various mixing operations. 3D DA has garnered attention recently, which also adopts similar methods such as scaling, rotation, and translation; Recent studies attempt to augment local structures of point clouds (li2020pointaugment, ; sheshappanavar2021patchaugment, ; kim2021point, ) and introduce concepts for object-level augmentation (chen2020pointmixup, ; lee2021regularization, ) and scene-level (xiao2022polarmix, ) in self-driving. However, conventional DAs are unsuitable for addressing the unique characteristics of wireless sensing data, as we discussed above, necessitating the development of customized approaches.
@@
Data augmentation for labeling efficiency. The mixing-based local DA can not only improve model performance in standard supervised learning, but also reduce labeling efforts via self-training (xiao2022polarmix, ). It enables unsupervised domain adaptation without explicit supervision. Its effectiveness has been studied in text analysis, computer vision (yuan2021simple, ; ghiasi2021simple, ; xiao2022polarmix, ), and speech processing (shorten2021text, ), improving the generalizability of models. It synthesizes silver standard labels generated from input data to facilitate learning. Similarly, in self-supervised learning, DA (swav2020caron, ; dino2021caron, ) even brings more performance improvements than algorithmic techniques like revising model architectures or algorithmic additions (simclr2020chen, ; mocov2021chen, ; byol2020grill, ; msn2022assran, ). This aspect is crucial for wireless perception because, unlike crawling images and text from the web, collecting sensing data with labels from the real physical world is especially expensive.
@@
Wireless sensing is a promising alternative to cameras for tracking humans, relying on the modalities of radio frequency (RF) signals, such as mmWave and WiFi, or acoustics. The applications range from gesture classification (hayashi2021radarnet, ; wang2016interacting, ), localization (zhang2021widar3, ; qian2018widar2, ; qian2017widar, ), motion detection and pose estimation (adib2013see, ; adib20143d, ; adib2015rf, ; zhao2018rf, ; zhao2018through, ; jiang2020towards, ), and fine-grained face reconstruction (xie2023mm3dface, ) using RF. In addition, acoustics detection leverages ubiquitous speakers and microphones with little bandwidth for hand gesture recognition (amaging2022, ; wang2020push, ; gupta2012soundwave, ; yang2023sequence, ) and hand tracking (nandakumar2016fingerio, ; wang2016device, ; mao2019rnn, ; liu2022acoustic, ; li2020fm, ).
@@
Wireless sensing datasets have become more accessible in recent research. Some focus primarily on a single task, mostly keypoint estimation or action recognition using mmWave (singh2019radhar, ; zhao2023cubelearn, ). MARS (marsdataset2021, ) is one of the pioneers, providing data for rehabilitation using mmWave. mRI (an2022mri, ) and MM-Fi (yang2024mm, ) are large-scale structures that contain around 160K frames. MiliPoint (cui2024milipoint, ) first includes all three main tasks in human tracking: user identification, action classification, and keypoint pose estimation, with a total of 49 different actions across 545k frames. These datasets facilitate research on wireless perception and make DA research possible in this domain by being able to benchmark comprehensively.
@@
Data augmentation in wireless sensing was mentioned in some work, typically tailored to their specific system, yet lacking a generic DA framework with comprehensive evaluations. For examples of global DA, MiliPoint (cui2024milipoint, ) employs a DA, named stack, which involves zero-padding and random resampling. Consequently, the augmented points essentially replicate the original ones with some subset and duplication. This would be a baseline in our experiments. (huang2019radiomodule, ; chen2023radiowavelet, ; clark2021training, ; blindsignal2019, ) use global DA like flipping, rotation, Gaussian noise, or wavelet but only to classify radio signal types, not applicable for sensing.  (acousticdataset2023, ) shifts the range profile along the range axis by a small scale, which will make data distorted beyond a certain augmentation scale, leading to decreased performance. Besides, (crosssenss2018, ) uses transfer learning to improve data efficiency across environments but it requires labeled data from the target domain.
@@
Recently, generative models have been explored to synthesize wireless data (zhao2023nerf2, ; chen2023rfgenesis, ), but are limited to certain tasks or data. (ahuja2021vid2doppler, ; 2023synthesizemili, ) simulates RF signals from videos of human actions, but it can only generate scenes from existing video data. (chen2023rfgenesis, ) overcomes this issue by using text prompts to synthesize the 3D mesh of humans first, and then simulate mmWave data from the 3D visuals using ray tracing and Diffussion models. Nevertheless, this only supports action-related tasks, because the text prompts can only describe body actions or hand gestures. For example, it can only synthesize data for pose estimation or action recognition tasks, but not for user identification, since it is intractable to accurately depict behavior traits solely by text prompts. In (zhao2023nerf2, ), NeRF is applied to simulate RF signals, but it can only aid simple static tasks of localization or 5G channel estimation rather than motion tracking, since the vanilla NeRF is for static synthesis; and it requires labeled data from the target scene to train the NeRF model. RF-Diffusion (chi2024rfdiff, ) is promising to generalize to tasks but it relies on conditional generation, which may face challenges with longer condition vectors, such as sequences of 90 or more coordinates; the paper only validated on action classification task. In summary, generative methods require significant effort to train an additional model and struggle to generalize to out-of-distribution data, while mixing-based DA is training-free and excels at enhancing data diversity.
3. Background: Wireless sensing and data formats
Figure 1. The common data processing pipeline for wireless perception.
Figure 2. The sparsity issue in wireless.
In this section, we first provide a brief background about traditional wireless sensing systems to reveal various challenges that lead to WixUp’s design.